A Zener diode regulator circuit


must


provide a constant voltage of 12V, across


a 250Ω load resistor. The circuit uses a 24V input DC source, a 12V Zener


diode and a 50Ω series resistance. Calculate the minimum acceptable power


rating of the Zener diode.



1
Expert's answer
2021-07-07T09:38:02-0400

Total current in the circuit

"I=\frac{24-12}{50}=\frac{12}{50}=0.24A\\
V_z=12V\\"

Current in 250ohms resistor

"I_R=\frac{12}{250}=0.048A"

Current through zener diode is

"I=I_z+I_R\\
I_z=0.24-0.048=0.192A\\"

Minimum acceptable power is

"P_z=I_z\cdot V_z\\
P_z=0.192\cdot 12=2.304W"